# Repetita Iuvant: Data Repetition Allows SGD to Learn High-Dimensional Multi-Index Functions

### v1

@@ 4.2 Learning hard functions through a hierarchical mechanism
The investigation of the hierarchical nature of SGD learning has attracted noticeable attention (Abbe et al., 2021, 2022, 2023; Dandi et al., 2023). In this paper, we portray a completely different picture in terms of computational efficiencies when data repetition is considered in the algorithmic SGD routine. One may wonder if there is a generalization of such a hierarchical learning mechanism to the present novel setting; we show that coupling between directions can hierarchically guide the learning process.
The Left panel in Fig. 3 shows an example of the so-called staircase functions (Abbe et al., 2021), i.e. $f^{\star}(\bm{z})=z_{1}+z_{1}\mathrm{He}_{3}(z_{2})$. While one-pass SGD needs to learn hierarchically first the direction $\bm{w}^{\star}_{1}$ in $T=O(d)$ and then $\bm{w}^{\star}_{2}$ in $T=O(d^{2})$, Algorithm 1 easily learns both in linear time. This observation could lead to infer that hierarchical learning mechanisms are not present when more realistic training scenarios are considered.
(a) $f^{\star}(\bm{z})=z_{1}+z_{1}\mathrm{He}_{3}(z_{2})$
(b) $f_{\mathrm{sign}}^{\star}(\bm{z})=\mathrm{sign}(z_{1}z_{2}z_{3}),\quad f^{\star}_{\mathrm{stair}}(\bm{z})=\mathrm{He}_{2}(z_{1})+\mathrm{sign}(z_{1}z_{2}z_{3})$
Figure 3: Hierarchical picture of learning – Evolution of the maximum Cosine Similarities with different target directions $\{\bm{w}^{\star}_{r}\}_{r\in[k]}$ are identified by different colors: $\bm{w}^{\star}_{1}$ (blue), $\bm{w}^{\star}_{2}$ (orange), $\bm{w}^{\star}_{3}$ (green) as a function of the normalized iteration time. (a): both the algorithms learn $\bm{w}^{\star}_{1}$ in linear time, but only EgD can also learn $\bm{w}^{\star}_{2}$ in $O(d)$ using staircase mechanism; SGD requires another $O(d^{2})$ steps to take advatage of the staircase and learn $\bm{w}^{\star}_{2}$. (b): EgD performance for the target function $f^{\star}_{\mathrm{stair}}$ (dots) and $f^{\star}_{\mathrm{sign}}$(squares); we plot in separate subplots the overlap with different directions to highlight the presence of hierarchical learning mechanism. Learning the first target direction $\bm{w}^{\star}_{1}$ triggers the hierarchical mechanism and EgD is able to weakly recover the full target subspace $\mathrm{Span}(\bm{w}^{\star}_{1},\bm{w}^{\star}_{2},\bm{w}^{\star}_{3})$ in $O(d\log d)$, while this does not happen when removing $\mathrm{He}_{2}(x_{1})$ from the target. See App. C for additional details.
To refute this, we illustrate in the right panel of Fig. 3 a scenario that precisely exemplifies the presence of hierarchical mechanisms within Algorithm 1. We run this algorithm on two different target functions, $f_{\mathrm{sign}}^{\star}(\bm{z})=\mathrm{sign}(z_{1}z_{2}z_{3})$ and $f^{\star}_{\mathrm{stair}}(\bm{z})=\mathrm{He}_{2}(z_{1})+\mathrm{sign}(z_{1}z_{2}z_{3})$. We show in App. A that $3-$sparse parities of the form $\mathrm{sign}(z_{1}z_{2}z_{3})$ are hard functions even for SGD with data repetition (Algorithm 1), and indeed they are not learned in (almost) linear time. On the other hand, for the function $f^{\star}_{\mathrm{stair}}$, our simulations predict that the first direction $\bm{w}^{\star}_{1}$ is learned in $T=O(d\log d)$ iterations (see the rightmost panel of Fig. 2). However, once the direction $\bm{w}^{\star}_{1}$ is learned, it can be used to obtain order-one correlation with $\{\bm{w}^{\star}_{2},\bm{w}^{\star}_{3}\}$ again in $T=O(d\log d)$ steps. While the latter example belongs to the class of “CSQ” staircase function depicted in Abbe et al. (2021), novel “SQ” hierarchical mechanisms arise in our framework, such as for instance the function $f_{\rm sq-stair}^{\star}(\bm{z})=\mathrm{He}_{4}(z_{1})+\mathrm{sign}(z_{1}z_{2}z_{3})$ (see Fig. 4).
@@
Figure 4: Novel hierarchical picture of learning – Evolution of the maximum Cosine Similarities attained by EgD with different target directions $\{\bm{w}^{\star}_{r}\}_{r\in[k]}$ are identified by different colors: $\bm{w}^{\star}_{1}$ (blue), $\bm{w}^{\star}_{2}$ (orange), $\bm{w}^{\star}_{3}$ (green) as a function of the normalized iteration time. The target function is $f^{\star}_{\rm sq-stair}(\bm{z})=\mathrm{He}_{4}(z_{1})+\mathrm{sign}(z_{1}z_{2}z_{3})$, that is an SQ-staircase, but not a CSQ-staircase. EgD learn $\bm{w}^{\star}_{1}$ in $O(d\log d)$ steps and use the information to learn the other two directions in another $O(d\log d)$ steps. SGD (not showed in the plot) cannot take advantage of the staircas mechanism since $\mathrm{He}_{4}(z_{1})$ and $\mathrm{sign}(z_{1}z_{2}z_{3})$ have information exponent $\ell=4$ and $\ell=3$ respectively. See App. C for additional details.
@@ 5 Conclusions
A natural avenue of future research would be to follow Damian et al. (2023) and use smoothing to make them even closer to SQ bounds in the more difficult setting where the generative exponent is higher than two (associated with a resctricted class of non-polynomial functions).
@@ Acknowledgements
The authors would like to thank Nicolas Flammarion, Bruno Loureiro, Nati Srebro, Emanuele Troiani, and Lenka Zdeborová for interesting discussions. This work was supported by the Swiss National Science Foundation under grant SNSF OperaGOST (grant number 200390).
Appendix A Proof of the Main Results
A.1 Proof of Theorem 2
We formalize here the proof sketch of Theorem 2. For simplicity, we assume that $a_{0}=1$. We track the dynamics of $\bm{w}_{t}$ through the following sufficient statistic:
$m_{t}:=\langle\bm{w}_{t},\bm{w}^{\star}\rangle.$ (14)
We assume in the following that $m_{0}>0$, so that $\epsilon=1$ in Assumption 3. We will discuss the actual choice of $\epsilon$ in the course of the proof.
We define the following quantities:
$\Phi(\bm{z})$ $=\nabla_{\bm{w}}\mathcal{L}\left(f(\bm{z};\tilde{\bm{w}}(\rho),a_{0}),y\right)$ (15)
$\Psi(x,x^{\star})$ $=h^{\star}(x^{\star})\sigma^{\prime}(x+\rho_{0}\sigma^{\prime}(x)h^{\star}(x^{\star}))x^{\star}$ (16)
$\phi(m)$ $=\mathbb{E}[\Psi(x,x^{\star})]\quad\text{for}\quad\dbinom{x}{x^{\star}}\sim\mathcal{N}\left(\bm{0},\begin{pmatrix}1&m\\ m&1\end{pmatrix}\right).$ (17)
Using these definitions, we have for the update equation (11)
@@ A.1 Proof of Theorem 2
$\bm{w}_{t+1}=\frac{\bm{w_{t}}-\gamma\Phi(\bm{z_{t}})}{\left\lVert\bm{w_{t}}-\gamma\Phi(\bm{z_{t}})\right\rVert}$
This places us under the framework of Ben Arous et al. [2021], replacing $\nabla\mathcal{L}_{N}$ by $\Phi$ (and noting that all conditions in their theorems relate to $\nabla\mathcal{L}_{N}$ instead of $\mathcal{L}_{N}$). We also define the constant $\kappa(\delta)$ such that

### v2

@@ 4.2 Learning hard functions through a hierarchical mechanism
The investigation of the hierarchical nature of SGD learning has attracted noticeable attention (Abbe et al., 2021, 2022, 2023; Dandi et al., 2024a). In this paper, we portray a completely different picture in terms of computational efficiencies when data repetition is considered in the algorithmic SGD routine. One may wonder if there is a generalization of such a hierarchical learning mechanism to the present novel setting. Indeed for optimal AMP algorithm, a diffrent picture emerges, (called Grand staircase in Troiani et al. (2024)). Here we show that coupling between directions can hierarchically guide the learning process as well with SGD.
@@
Figure 4: Hierarchical learning – Evolution of the maximum Cosine Similarities attained by EgD with different target directions $\{\mathbf{w}^{\star}_{r}\}_{r\in[k]}$ are identified by different colors: $\mathbf{w}^{\star}_{1}$ (blue), $\mathbf{w}^{\star}_{2}$ (orange), $\mathbf{w}^{\star}_{3}$ (green) as a function of the normalized iteration time. The target function is $f^{\star}_{\mathrm{sq-stair}}(\mathbf{z})=\mathrm{He}_{4}(z_{1})+\mathrm{sign}(z_{1}z_{2}z_{3})$, that is an SQ-staircase (or grand staircase), not a CSQ-staircase. EgD learn $\mathbf{w}^{\star}_{1}$ in $O(d\log d)$ steps and use the information to learn the other two directions in another $O(d\log d)$ steps. SGD (not showed in the plot) cannot take advantage of the staircase mechanism since $\mathrm{He}_{4}(z_{1})$ and $\mathrm{sign}(z_{1}z_{2}z_{3})$ have information exponent $\ell=4$ and $\ell=3$ respectively. (See App. E).
The Left panel in Fig. 3 shows an example of the so-called staircase functions (Abbe et al., 2021), i.e. $f^{\star}(\mathbf{z})=z_{1}+z_{1}\mathrm{He}_{3}(z_{2})$. While one-pass SGD needs to learn hierarchically first the direction $\mathbf{w}^{\star}_{1}$ in $T=O(d)$ and then $\mathbf{w}^{\star}_{2}$ in $T=O(d^{2})$, Algorithm 1 easily learns both in linear time. This observation could lead to infer that hierarchical learning mechanisms are not present when more realistic training scenarios are considered.
To refute this, we illustrate in the right panel of Fig. 3 a scenario that precisely exemplifies the presence of hierarchical mechanisms within Algorithm 1. We run this algorithm on two different target functions, $f_{\mathrm{sign}}^{\star}(\mathbf{z})=\mathrm{sign}(z_{1}z_{2}z_{3})$ and $f^{\star}_{\mathrm{stair}}(\mathbf{z})=\mathrm{He}_{2}(z_{1})+\mathrm{sign}(z_{1}z_{2}z_{3})$. We show in App. B that $3-$sparse parities of the form $\mathrm{sign}(z_{1}z_{2}z_{3})$ are hard functions even for SGD with data repetition (Algorithm 1), and indeed they are not learned in (almost) linear time. On the other hand, for the function $f^{\star}_{\mathrm{stair}}$, our simulations predict that the first direction $\mathbf{w}^{\star}_{1}$ is learned in $T=O(d\log d)$ iterations (see the rightmost panel of Fig. 2). However, once the direction $\mathbf{w}^{\star}_{1}$ is learned, it can be used to obtain order-one correlation with $\{\mathbf{w}^{\star}_{2},\mathbf{w}^{\star}_{3}\}$ again in $T=O(d\log d)$ steps. While the latter example belongs to the class of “CSQ” staircase function depicted in Abbe et al. (2021), novel “SQ” hierarchical mechanisms arise in our framework, such as for instance the function $f_{\mathrm{sq-stair}}^{\star}(\mathbf{z})=\mathrm{He}_{4}(z_{1})+\mathrm{sign}(z_{1}z_{2}z_{3})$ (see Fig. 4).
The staircase picture in terms of generative exponent was discussed for AMP-type algorithm in Troiani et al. (2024). Our study indicate that this new class (they called grand staircase) is not restricted to AMP: We anticipate that grand staircase functions can be efficiently learned by two-layer networks when data reusing is permitted, whereas only standard staircase ones emerge under single-pass SGD. Additional recent results by (Joshi et al., 2024) align with this perspective.
@@ Acknowledgements
Upon completion of this work, we became aware of the interesting paper from (Lee et al., 2024) showing strong recovery with similar sample complexity, also by exploiting the reuse of training data. Our work was conducted independently and simultaneously. The authors would like to thank Nicolas Flammarion, Bruno Loureiro, Nati Srebro, Emanuele Troiani, and Lenka Zdeborová for interesting discussions. This work was supported by the Swiss National Science Foundation under grant SNSF OperaGOST (grant number 200390).
Appendix A Theoretical framework
A.1 Difference inequalities for vector-valued processes
We begin with a lemma that generalizes [Ben Arous et al., 2021, Proposition 4.1]. Let $\Phi:\mathbb{R}^{d}\to\mathbb{R}^{d}$ be a random function, and consider the process $(\mathbf{u}_{t})_{t\geq 0}\in\mathbb{S}^{d-1}$ given by
$\mathbf{u}_{t+1}=\frac{\mathbf{u}_{t}+\gamma\Phi(\mathbf{u}_{t})}{\left\lVert\mathbf{u}_{t}+\gamma\Phi(\mathbf{u}_{t})\right\rVert}.$
We make the following assumptions:
Assumption 6.
The initial value of $\mathbf{u_{0}}\in\mathbb{S}^{d-1}$ is approximately isotropic, in the sense that there exists a constant $c>0$ such that
$\mathbb{E}\left[\mathbf{u_{0}}\mathbf{u_{0}}^{\top}\right]\succ\frac{c}{d}I$
Assumption 7.
There exist constants $C,\iota>0$ such that
$\sup_{\mathbf{u}\in\mathbb{S}^{d-1}}\,\left\lVert\mathbb{E}\left[\Phi(\mathbf{u})\Phi(\mathbf{u})^{\top}\right]\right\rVert$ $\leq C$ (15)
@@ A.1 Difference inequalities for vector-valued processes
$\sup_{\mathbf{u}\in\mathbb{S}^{d-1}}\,\mathbb{E}\left[\left\lVert\Phi(\mathbf{u})\right\rVert^{4+\iota}\right]$ $\leq CN^{\frac{4+\iota}{2}}$ (16)
Further, we have $\langle\mathbf{u},\Phi(\mathbf{u})\rangle=0$ a.s. for any $\mathbf{u}\in\mathbb{S}^{d-1}$.
We fix a vector $\mathbf{v}$ and consider the random process $m_{t}(\mathbf{v}):=\langle\mathbf{u}_{t},\mathbf{v}\rangle$. For any $s\geq 0$, we define the stopping times
$\tau_{\theta}^{+}(\mathbf{v})=\inf\{t\geq 0:m_{t}(\mathbf{v})\geq\theta\}\quad\text{and}\quad\tau_{\theta}^{-}(\mathbf{v})=\inf\{t\geq 0:m_{t}(\mathbf{v})\leq\theta\}$
Proposition 1.
Fix an $\alpha>0$. There exist constants $c(\delta,\alpha),\kappa(\delta)>0$ such that if
$\gamma\leq c(\delta,\alpha)d^{-1}\quad\text{and}\quad T\leq c(\delta,\alpha)\gamma^{-2}d^{-1},$ (17)
then, with probability $1-\delta$, the following holds. The initialization $m_{0}(\mathbf{v})$ satisfies
$|m_{0}(\mathbf{v})|\geq\frac{\kappa}{\sqrt{d}},$ (18)
and:
  • if $m_{0}(v)>0$, for any $t\leq T\wedge\tau_{0}^{-}(\mathbf{v})$,
    $(1-\alpha)m_{0}(\mathbf{v})+\gamma\sum_{s=0}^{t-1}\langle\Psi^{+}(\mathbf{u}_{s}),\mathbf{v}\rangle\leq m_{t}(\mathbf{v})\leq(1+\alpha)m_{0}(\mathbf{v})+\gamma\sum_{s=0}^{t-1}\langle\Psi(\mathbf{u}_{s}),\mathbf{v}\rangle,$ (19)
  • if $m_{0}(v)<0$, for any $t\leq T\wedge\tau_{0}^{+}(\mathbf{v})$,
    $(1+\alpha)m_{0}(\mathbf{v})+\gamma\sum_{s=0}^{t-1}\langle\Psi(\mathbf{u}_{s}),\mathbf{v}\rangle\leq m_{t}(\mathbf{v})\leq(1-\alpha)m_{0}(\mathbf{v})+\gamma\sum_{s=0}^{t-1}\langle\Psi^{-}(\mathbf{u}_{s}),\mathbf{v}\rangle,$ (20)
where $\Psi,\Psi^{\pm}$ is a deterministic function equal to
$\Psi(\mathbf{u})=\mathbb{E}\left[\Phi(\mathbf{u})\right]\quad\text{and}\quad\Psi^{\pm}(\mathbf{u})=\Psi(\mathbf{u})\mp\gamma\mathbb{E}\left[\left\lVert\Phi(\mathbf{u})\right\rVert^{2}\right]\mathbf{u}.$
Proof.
We assume for simplicity that $m_{0}(\mathbf{v})>0$; the case where $m_{0}(\mathbf{v})>0$ is treated identically. From Assumption 6, by a simple Markov inequality, there exists a constant $\kappa(\delta)>0$ such that